# Model Compression in Practice: Lessons Learned from Practitioners Creating On-device Machine Learning Experiences

### v1

@@ Abstract.
Efficient machine learning, machine learning model compression, on-device machine learning, interview study, interactive systems, design directions
$c$s: Human-centered computing Interactive systems and toolsccs: Computing methodologies Machine learningccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Most modern machine learning (ML) models in production today occupy cloud servers of exceedingly more computational capacity and power than the device in your pocket. Yet for delivering intelligent user experiences, there is good reason to shift ML models onto our everyday personal computing devices.
On-device ML is the practice of storing, training, and running ML models on an individual’s device, such as a smartphone, tablet, or wearable. However, as today’s state-of-the-art models grow larger and larger in size (e.g., into the billions of parameters) (Villalobos et al., 2022; Giattino et al., 2022; Zhao et al., 2023; Stanford, 2023), efficiency remains the biggest barrier to on-device ML (Apple, 2023a; Developers, 2022). An arbitrary ML model placed on a mobile device can easily consume every available resource of the device, whether it be compute, memory, or battery. Creating efficient, on-device models brings new challenges to the ML development process.
In this paper we argue that research in efficient machine learning over the past decade has matured to the degree to where practitioners have an (albeit rapidly evolving) set of techniques to make on-device ML a reality. The key idea is to shrink, optimize, and compress models, while simultaneously maintaining their accuracy. To achieve this, practitioners develop strategies for how to best apply model compression techniques to minimize the amount of computational resources needed. Product designers, tool-builders, and ML practitioners today should be considering the benefits of on-device ML for their users:
First, on-device models can be an enormous win for personal privacy. While the industry standard practice today is to send a user’s (encrypted) data over a network to a commercially-owned server for ML inference, on-device ML cuts out the need for a server such that a user’s personal data never leaves their device. Beyond inference, pre-trained models can be fine-tuned on-device to adapt and personalize to an individual user’s preferences, while keeping those preferences private and local (Konečnỳ et al., 2016; Hu et al., 2020; Apple, 2023b).
@@
Lastly, removing the reliance on servers for ML inference has economic and environmental impact. On-device ML cuts out the cost of server, which may help individuals, non-profits, or smaller tech firms that previously could not afford server upkeep to now deliver ML-powered features to their users. By reducing society’s reliance on external servers, on-device ML may help reduce the high carbon footprint of the cloud (Monserrate, 2022).
In this paper, we seek to help the broader community adopt on-device ML by filling a crucial gap: while efficient ML research and algorithmic advances have progressed (Zhao et al., 2022; Gu et al., 2021; Zamzam et al., 2019; Dhar et al., 2021; Sehgal and Kehtarnavaz, 2019), there currently exists little pragmatic guidance (NVIDIA, 2023) in literature, online, in books, or otherwise, for newcomers wanting to create on-device ML users experiences. To curate pragmatic guidance, we interviewed 30 industry ML practitioners (research scientists, ML engineers, and product managers) who are uniquely experienced with designing, developing, and deploying on-device ML at scale. We capture the hard-won knowledge, insights, and tricks-of-the-trade that are traditionally omitted from technical ML publications to illustrate the current state-of-the-art. We also profile how industry practitioners adapt cutting-edge ML research into intelligent experiences running on peoples’ everyday devices. We contribute:
  • Tacit knowledge that 30 expert practitioners have developed around the design process, trade-offs, and strategies that go into deploying efficient on-device ML. Many decisions around optimization and compression strategy stem directly from user experience and product design.
  • A characterization of the key challenges practitioners face when creating efficient models. Example challenges include the tension between optimizing performance against model accuracy, and the necessity to work with hardware-level details.
@@
We conclude by discussing where the broader human-centered AI/ML research, design, and practitioner communities can engage in efficient, on-device ML. We hope the results from this deep dive into a nascent area of ML user experience design helps spotlight its interdisciplinary importance and inspires others to contribute.
@@ 2. Related Work in Human-Centered ML and Model Optimization
Over the past decade, interview studies from human-computer interaction (HCI) have played an important role in capturing how AI/ML plays out in real-world practice (Amershi et al., 2019a, b; Holstein et al., 2019; Sambasivan et al., 2021; Liao et al., 2020; Bhatt et al., 2020; Yang et al., 2018). Interviews with practitioners are uniquely suited to capture the kinds of process-oriented insight, stories, and tricks-of-the-trade that are traditionally omitted from technical ML publications. Interviews also crucially give the broader public access to learn from AI/ML work that otherwise happens behind closed doors. In this paper, we aim to share new ideas for designing powerful on-device ML experiences, based on extensive product experience that is currently held by few. To situate our work in the broader landscape of HCI + ML research, we cover other interviews studies within diverse disciplines of AI/ML research, including user experience design, best practices for development, and interactive systems and tools.
2.1. User Experience Design for AI/ML
Notably, much of what the research community knows today about designing effective user experiences (UX) for AI/ML has been taught through the hard-won experience of real designers. For example, Yang et al. interviewed 13 UX practitioners with 4+ years of experience working on AI/ML products, to capture designers’ strategies for working with AI/ML (Yang et al., 2018). Zdanowska et al. interviewed both product teams and UX designers to explore how design collaboration happens during AI/ML product development (Zdanowska and Taylor, 2022). Windl et al. interviewed another 20 interaction designers to learn how designers adapt their their innovation process for AI/ML experiences (Windl et al., 2022). Again with experienced UX + AI/ML designers, Yildirim et al. conducted workshops around AI/ML innovation and found that designers are more impactful when allowed to design entire services—rather than fitting AI/ML into specific user-interfaces for existing products (Yildirim et al., 2022). Meanwhile, leading technology companies have distilled design wisdom from their own product teams into AI/ML design guidelines as public, educational resources (Wright et al., 2020). Examples include Apple’s Human Interface Guidelines on Machine Learning (gui, 2019a), Microsoft’s Human-AI Guidelines (Amershi et al., 2019b), Google’s People + AI Guidebook (gui, 2019b), and IBM’s Design for AI resources (des, 2019).
This body of work emphasizes that designing with machine learning requires new approaches from designers (Yang, 2018; Amershi et al., 2019b; Windl et al., 2022; Yildirim et al., 2022; gui, 2019b; des, 2019). Our paper contributes to this conversation by illustrating how power and performance issues of moving ML models on-device create real, tangible UX constraints for designers (for example, see Table 2)—and offering concrete design strategies for creating effective user experiences around those constraints (for example, see Section 5.4).
2.2. Profiling Best Practices for Building AI/ML
@@
Regarding the engineering of AI/ML development, practitioner stories inform best-practices and new research agendas. Amershi et al. interview and survey AI/ML practitioners at Microsoft to gather best practices for production ML development and outline core differences between ML development and conventional software engineering (Amershi et al., 2019a). Researchers have studied how AI/ML teams collaborate across diverse technical roles (Passi and Jackson, 2018; Zhang et al., 2020; Piorkowski et al., 2021), or adopt new technology like AutoML (Wang et al., 2019, 2021b, 2021a). Sambasivan et al. used interviews to profile organizational struggles between balancing modeling work with data quality work (Sambasivan et al., 2021). Holstein et al. found key mismatches between academic AI/ML fairness concepts and the kinds of real fairness issues product teams grapple with (Holstein et al., 2019). Many other aspects of AI fairness, accountability, transparency, and ethics (FATE) have been found to be fundamentally sociotechnical issues where design decisions and process have enormous impact (Hong et al., 2020; Madaio et al., 2020; Modarres et al., 2018; Delgado et al., 2021; Madaio et al., 2022; Robertson et al., 2021; Rakova et al., 2021; Hopkins and Booth, 2021).
From a practical perspective within the ML community, existing work highlights the hidden technical debt within ML systems that incur massive amounts of engineering maintenance (Sculley et al., 2015). Some these ideas can be traced back before the proliferation of ML (Hand, 1998). More recently, related to our work is a recent technical survey on efficient methods for natural language processing (NLP) (Treviso et al., 2022) that aims to inspire new technical methods for efficiency in the age of ever-scaling language models; however, its primary focus is to only organize existing techniques specific to NLP applications. In our work, we profile on-device ML holistically, including its impact on product design, organization-level coordination, and best-practices for ML teams.
2.3. Interactive Systems for AI/ML Development
Extensive prior work in HCI, visualization, and ML has focused on interactive systems and tools to help ML practitioners overcome key challenges in their work. Examples include tools for: neural network architecture visualization and debugging (Wongsuphasawat et al., 2017), visualization for large ML dataset exploration (Google, 2021b, a) and data iteration (Hohman et al., 2020), model evaluation for classification tasks (Görtler et al., 2022; Amershi et al., 2015; Ren et al., 2016), visual analytics for deep learning interpretability (Hohman et al., 2018), fairness evaluation for finding disproportionately underperforming subpopulations of data (Bird et al., 2020; Cabrera et al., 2019), reliable error analysis and evaluation for NLP models (Wu et al., 2019; Ribeiro et al., 2020; Ribeiro and Lundberg, 2022), interactive visualization for counterfactual generation (Wu et al., 2021; Wexler et al., 2019), and interaction techniques to enhance ML and data science environments (Patel et al., 2008), such as computational notebooks (Kery et al., 2020) and other composable, shareable interfaces (Bäuerle et al., 2022). Today, on-device and efficient machine learning has little tooling available outside of code libraries to run compression algorithms (see applications within popular ML toolkits in Section 2.6). Our results from interviewing 30 experts provide strong designs recommendations for interactive tooling opportunities that would make on-device ML more achievable for everyone.
2.4. On-device Inference versus On-device Training
It is worthwhile to note a distinction between on-device inference and on-device training. On-device ML inference refers to running an ML model on any device with limited compute resources, such as limited compute, memory, or power. These models are pretrained, where learning happened offline on a different device or server. On-device ML training refers to either training a model from scratch, or fine-tuning a model for a few iterations, on a user’s device.
@@ 2.4. On-device Inference versus On-device Training
While work in on-device ML encompasses both paradigms, in our paper we focus on the far more common use case of model inference (deploying pretrained models on a user’s device), and leave on-device training for future work. Although on-device inference and training share many of the same core challenges around efficiency, training on-device models is generally considered harder, since training usually requires far more resources (Zhou et al., 2019). Thus, we do not attempt to cover the added complexities of on-device training in this work. For a survey on the current research and challenges around on-device learning instead, see: (Dhar et al., 2021; Lim et al., 2020; Zhou et al., 2019; Murshed et al., 2021).
2.5. Edge Computing
In recent literature, the challenges of on-device ML are often discussed in the context of edge computing (Zhou et al., 2019; Murshed et al., 2021; Chen and Ran, 2019). Edge computing typically refers to small embedded hardware, wearables, or internet-of-things (IoT) devices where at least some computation is performed on-device (the “edges” of a network) rather than by a central server (Murshed et al., 2021). In our research, we interview ML practitioners deploying to a variety of consumer devices, including edge hardware. We do not distinguish between device type except for where practices slightly differ for edge hardware, as in Section 5.4.4.
2.6. Existing Compression Resources
@@
Beyond academic surveys detailing different compression techniques (Deng et al., 2020; Choudhary et al., 2020; Cheng et al., 2018), most practical guidance for model optimization is found within online tutorials and documentation for popular ML libraries. Examples include TensorFlow’s model optimization toolkit and blog post on quantization-aware training method (TensorFlow, 2020, 2018); PyTorch’s experimental support for quantization (PyTorch, 2018), sparsity (PyTorch, 2019), and it’s accompanying examples (PyTorch, 2023); Google’s quantization extension to Keras called QKeras (Google, 2019); Microsoft’s Neural Network Intelligence package and tool (Microsoft, 2021); Intel’s Neural Compressor library (Intel, 2020); and Apple’s DNIKit (Welsh, 2023). Other examples that target specific hardware include speeding up inference on FPGAs (Fahim et al., 2021) and compressing Core ML models to run on Apple platforms (Apple, 2023a). In terms of other community efforts, the appropriately named TinyML community has sprung up around this topic, which has published a book (Warden and Situnayake, 2019) and hosts community events and meetups for like-minded practitioners.
2.7. Compressing Large Language Models and Foundation Models
Within the past couple of years, model compression has garnered intense interest due to the public proliferation of large language models (LLMs) and foundation models (Bommasani et al., 2021). LLMs have moved beyond the research community and captured public interest and concern (Solaiman et al., 2023; Movva et al., 2023), as they have demonstrate new ways to interact with computers via natural language. The catch is that these models, by definition, are the largest in the world (e.g., billions of parameters (Giattino et al., 2022; Villalobos et al., 2022)). For all the reasons already mentioned, there is immense interest and activity of moving LLMs on-device—and progress is moving fast. The number of empirical works for compressing LLMs from this year alone would be difficult to track. However, we note that the same overall approaches for efficient machine learning that we discuss in this work (detailed in Section 3) are the same techniques being applied to LLMs. For interested readers in learning about the latest results for compressing LLMs, we refer to the follow surveys: (Treviso et al., 2023; Zhou et al., 2023; Xu and McAuley, 2023).
3. A Primer on Model Compression Techniques

### v2

@@ Abstract.
Efficient machine learning, model compression, on-device machine learning, interview study, interactive systems, design directions
journalyear: 2024copyright: rightsretainedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642109isbn: 979-8-4007-0330-0/24/05submissionid: 3200ccs: Human-centered computing Interactive systems and toolsccs: Computing methodologies Machine learningccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Most modern machine learning (ML) models in production today occupy cloud servers of exceedingly more computational capacity and power than the device in your pocket. Yet for delivering intelligent user experiences, there is good reason to move ML models onto personal computing devices people use every day.
On-device ML is the practice of storing, training, and running ML models on an individual’s device, such as a smartphone, tablet, or wearable. However, as today’s state-of-the-art models grow larger and larger in size (e.g., into the billions of parameters (Villalobos et al., 2022; Giattino et al., 2022; Zhao et al., 2023; Stanford, 2023)), efficiency remains the biggest barrier to on-device ML (Apple, 2023a; Developers, 2022). An arbitrary ML model placed on a mobile device can easily consume every available resource of the device, whether it be compute, memory, or battery. Creating efficient, on-device models brings new challenges to the ML development process.
In this paper, we argue that research in efficient machine learning over the past decade has matured to the degree to where practitioners have an (albeit rapidly evolving) set of techniques to make on-device ML a reality. The key idea is to shrink, optimize, and compress models, while simultaneously maintaining their accuracy. To achieve this, practitioners develop strategies for how to best apply model compression techniques to minimize the amount of computational resources needed. Product designers, tool-builders, and ML practitioners today should be considering the benefits of on-device ML for their users:
First, on-device ML can be an enormous win for personal privacy. While standard practice today is to send a user’s (encrypted) data over a network to servers for ML inference, on-device ML cuts out this dependency such that a user’s personal data never leaves their device. Beyond inference, pre-trained models can be fine-tuned on-device to adapt to an individual user’s preferences, while keeping those preferences private and local (Konečnỳ et al., 2016; Hu et al., 2020; Apple, 2023b).
@@
Lastly, removing the reliance on servers for ML inference has economic and environmental impact. On-device ML cuts out the cost of a server, which may help individuals, non-profits, or smaller tech firms that previously could not afford server upkeep to now deliver ML-powered features to their users. By reducing society’s reliance on external servers, on-device ML may help reduce the carbon footprint of the cloud (Monserrate, 2022).
In this paper, we seek to advocate for on-device ML by filling a crucial gap: while efficient ML research and algorithmic advances have progressed tremendously (Menghani, 2023; Zhao et al., 2022; Gu et al., 2021; Zamzam et al., 2019; Dhar et al., 2021b; Sehgal and Kehtarnavaz, 2019), there currently exists little pragmatic guidance (NVIDIA, 2023) in literature, online, in books, or otherwise, for people wanting to create on-device ML user experiences. On-device ML requires both clever algorithmic and user experience design, making this a fundamentally interdisciplinary problem that has received limited attention outside of ML venues. Our work addresses: RQ: How should a broader audience of HCI and ML practitioners today optimize powerful models to design on-device, ML user experiences?
To curate pragmatic guidance, we interviewed 30 expert industry practitioners who are uniquely experienced with designing, developing, and deploying on-device ML at scale. We capture the hard-won knowledge, insights, and tricks-of-the-trade that are traditionally omitted from technical ML publications. We draw connections between the design of ML user experiences and the compression strategies experts use to match design goals. We contribute:
  • Tacit knowledge that 30 expert practitioners have developed around the design processes, trade-offs, and strategies that go into deploying efficient on-device ML. Many decisions around optimization and compression strategy stem directly from user experience and product design.
  • A characterization of the key challenges practitioners face when creating efficient models. Examples include the tension between optimizing performance against accuracy, and the necessity to work with hardware-level details.
@@
We conclude by discussing where the broader HCI + AI/ML research, design, and practitioner communities can engage in efficient, on-device ML. We hope the results from this deep dive into a nascent area of ML user experience design helps spotlight its interdisciplinary importance and inspires others to contribute.
@@ 2. Related Work in Human-Centered ML and Model Optimization
Our work attempts to bridge the Machine Learning, Ubiquitous Computing, Mobile Computing, and Human–Computer Interaction (HCI) communities, to build common language around an ML topic with broad, intersectional impact. Our biggest challenge is in situating this research where no direct prior work exists. In this section, we anchor our work in two main threads of HCI research. First, overlapping with the HCI community, the Ubicomp and Mobile computing communities have long used model compression for on-device ML to make new user experiences possible (Dhar et al., 2021a). Our work complements this literature by consolidating together strategies across many use cases, with the aim of broadening the audience of practitioners who can contribute to creating on-device ML systems (Section 2.1). Second, our research continues a line of HCI work synthesizing lessons from real-world ML practice (Amershi et al., 2019a, b; Holstein et al., 2019; Sambasivan et al., 2021; Liao et al., 2020; Bhatt et al., 2020; Yang et al., 2018), and we extend prior work into a previously unconsidered area of ML (Section 2.2). We situate our work in both of these directions, and end by covering related work on model efficiency for on-device ML itself (Section 2.3).
2.1. Ubiquitous Computing: Deploying ML on Mobile & Edge Devices
Challenges of on-device ML often occur in the context of mobile or edge computing (Zhou et al., 2019; Murshed et al., 2021; Chen and Ran, 2019). Edge computing refers to small embedded hardware, wearables, or internet-of-things (IoT) devices where at least some computation is performed on-device (the “edges” of a network) rather than by a central server (Murshed et al., 2021). Example uses of compressed neural networks include facial recognition on mobile devices (Zhen et al., 2021), activity recognition on cameras (Mishra et al., 2020), gesture recognition on smartwatches (Xu et al., 2022), and respiratory monitoring on phones and smartwatches (Chauhan et al., 2018; Dai et al., 2021; Liaqat et al., 2019). Since efficiency is critical for edge devices, the Ubiquitous computing, Mobile computing, and IoT research communities have contributed model compression advances around dynamic models (Mishra and Gupta, 2023; Liu et al., 2021; Wang et al., 2023), structured sparsity (Liberis and Lane, 2023), and on-device training (Yao et al., 2021; Jiang et al., 2021; Li et al., 2023). Our work differs from prior literature by not focusing on any single application or technique, and instead synthesizing practitioner strategies for model compression that can be used across use cases and hardware types. In our research, we interviewed ML practitioners deploying to a variety of consumer devices, including edge hardware. We distinguish between device type only where advice differs for small embedded hardware, for example in Section 5.4.4.
2.2. Human-Computer Interaction: Studying AI/ML Practitioners
@@
Over the past decade, interview studies from HCI have played an important role in giving the public access to learn from AI/ML work that otherwise happens behind closed doors (Amershi et al., 2019a, b; Holstein et al., 2019; Sambasivan et al., 2021; Liao et al., 2020; Bhatt et al., 2020; Yang et al., 2018). Interviews with practitioners are uniquely suited to capture the kinds of process-oriented insight, stories, and tricks-of-the-trade that are traditionally omitted from technical ML publications. For instance, Amershi et al. (2019a) interview and survey AI/ML practitioners at Microsoft to gather best practices for production ML development. Researchers have studied how AI/ML teams collaborate across diverse technical roles (Passi and Jackson, 2018; Zhang et al., 2020; Piorkowski et al., 2021), or adopt new technology like AutoML (Wang et al., 2019, 2021b, 2021a). Sambasivan et al. (2021) used interviews to profile organizational struggles between balancing modeling work with data quality work. Holstein et al. (2019) found key mismatches between academic AI/ML fairness concepts and the kinds of real fairness issues product teams grapple with. Many other aspects of AI fairness, accountability, transparency, and ethics have been found to be practitioner-driven, where design decisions and process have enormous impact (Hong et al., 2020; Madaio et al., 2020; Modarres et al., 2018; Delgado et al., 2021; Madaio et al., 2022; Robertson et al., 2021; Rakova et al., 2021; Hopkins and Booth, 2021). Like this prior work, we examine the production and practice side of on-device ML efficiency to illustrate the connection between ML compression choices and their impact on holistic ML user-experience design.
2.2.1. User Experience Design for AI/ML
Much of what the research community knows today about designing effective user experiences (UX) for AI/ML comes from interviews with seasoned product designers (Yang et al., 2018; Zdanowska and Taylor, 2022; Windl et al., 2022; Yildirim et al., 2022). Leading technology companies have distilled design wisdom from their own product teams into AI/ML design guidelines as public, educational resources (Wright et al., 2020). Examples include Apple’s Human Interface Guidelines on Machine Learning (gui, 2019a), Microsoft’s Human-AI Guidelines (Amershi et al., 2019b), Google’s People + AI Guidebook (gui, 2019b), and IBM’s Design for AI resources (des, 2019). This body of work emphasizes that designing with machine learning requires new approaches from designers (Yang, 2018; Amershi et al., 2019b; Windl et al., 2022; Yildirim et al., 2022; gui, 2019b; des, 2019). Our paper contributes to this conversation by illustrating how power and performance issues of moving ML models on-device create real, tangible UX constraints for designers (for example, see Table 2), and offers concrete design strategies for creating effective user experiences around those constraints (for example, see Section 5.4).
2.3. Machine Learning: On-Device ML & Compression Techniques
The literature on efficient machine learning is broader than the scope of this paper, and for a comprehensive look we refer readers to the excellent survey papers cited here and below (Menghani, 2023; Deng et al., 2020; Choudhary et al., 2020; Cheng et al., 2018; Treviso et al., 2023). For neural networks, Menghani (2023) breaks efficiency into 5 areas: (1) compression techniques, (2) learning techniques, (3) automation, (4) efficient architectures, and (5) infrastructure & hardware. Our interviews with practitioners contained more discussion of (1) and (5), and some discussion of (2) and (4). We do not claim that this interview study is a comprehensive look at efficient machine learning. Nonetheless, our work adds a fresh perspective to existing efficient ML literature: instead of detailing specific optimization techniques, here we profile higher-level strategies for how practitioners put techniques into practice to enable user-experience goals.
2.3.1. On-device Inference vs. On-device Training
@@ 2.3. Machine Learning: On-Device ML & Compression Techniques
It is worthwhile to distinguish between on-device inference and on-device training. On-device inference refers to running an ML model on new input to get a prediction. These models are typically pretrained on a server, then delivered to a device. On-device training refers to either training a model from scratch or fine-tuning a model on a user’s device. While work in on-device ML encompasses both paradigms, in our paper we focus on the (currently) more common use case of on-device inference and leave on-device training for future work. Training on-device is generally considered harder, since training usually requires far more resources (Zhou et al., 2019). For a survey on the current challenges around on-device learning, see: (Dhar et al., 2021b; Lim et al., 2020; Zhou et al., 2019; Murshed et al., 2021).
2.3.2. On-device Large Language Models and Foundation Models
In recent years, large language models (LLMs) and other generalized foundation models have raised the magnitude of size we expect from ML models (Solaiman et al., 2023; Movva et al., 2023; Giattino et al., 2022; Villalobos et al., 2022). As models get bigger, so do the stakes for model compression and on-device efficiency (Bommasani et al., 2021). Beyond speeding up inference (Alizadeh et al., 2023), efficiency methods for LLMs also aim to lower their enormous pretraining and fine-tuning costs. Examples include low-rank adaption methods (e.g., LRPD (Zhao et al., 2016) and LoRA (Hu et al., 2022)), and parameter-efficient fine-tuning methods (Ding et al., 2023, 2022; Fu et al., 2023; Lialin et al., 2023). The number of empirical works for compressing LLMs from 2023 alone would be difficult to capture, therefore we refer interested readers to the following surveys: (Treviso et al., 2023; Zhou et al., 2023; Xu and McAuley, 2023). We note that many of the techniques for compression that we discuss in Section 3 are the same key ideas being applied to LLMs and foundation models.
2.3.3. Publicly Available Compression Resources
@@
While most compression techniques originate in academic work (Deng et al., 2020; Choudhary et al., 2020; Cheng et al., 2018), many resources for practitioners can be found within web tutorials and ML toolkits. Examples include TensorFlow’s quantization-aware training method (TensorFlow, 2020, 2018); PyTorch’s experimental support for quantization (PyTorch, 2018), sparsity (PyTorch, 2019), and it’s accompanying examples (PyTorch, 2023); Google’s quantization extension to Keras called QKeras (Google, 2019); Microsoft’s Neural Network Intelligence package and tool (Microsoft, 2021); Intel’s Neural Compressor library (Intel, 2020); and Apple’s MLX framework (Hannun et al., 2023) and DNIKit (Welsh et al., 2023). Other examples that target specific hardware include speeding up inference on FPGAs (Fahim et al., 2021) and compressing Core ML models to run on Apple platforms (Apple, 2023a). In terms of other community efforts, the appropriately named TinyML community has published a book (Warden and Situnayake, 2019) and hosts community events and meetups.
3. A Primer on Machine Learning Compression Techniques